I have 2 hectares of land. How long should my fence be in order to enclose it ? in meters.
Answer by Theo(13342)

1 hectare is equal to 10,000 square meters.
you have 2 hectares so you have 20,000 square meters of land.
the fencing required for this land will depend on the shape of the area.

if it is in the shape of a square, then each side is equal to sqrt(20,000) = 141.4213562 meters and the perimeter is 4 * that which is equal to 565.6854249 meters.
you would need 565.6854249 meters of fencing to surround it.

if it is in the shape of a rectangle and the short side of the rectangle is equal to 100 meters, then the long side of the rectangle is equal to 20,000 / 100 = 200 meters.
the area is 100 * 200 = 20,000 square meters
the perimeter is equal to 2 * 100 + 2 * 200 = 200 + 400 = 600 meters.
you would need 600 meters of fencing to surround it.

if it was in the shape of a circle, then the area is 20,000 square feet and you can solve for the radius by using the followng formula:
pi * r^2 = 20,000
solve for r to get r = sqrt(20,000/pi) = 79.78845608 meters.
the perimeter of a circle is equal to 2 * pi * r.
this makes the perimeter of the circular land equal to 2 * pi * 79.78845608 = 501.3256549 meters.

the area is the same regardless of the shape, but the perimeter is different depending on the shape.

in order to find out hos much fencing you require to enclose the land, you need to know what the shape of the land is.